OpenStudy (anonymous):

what is 84 feet per hour converted into inches per minute?

OpenStudy (tkhunny):

Multiply by 1 - as many times as necessary. \(\dfrac{84\;ft}{hr}\cdot\dfrac{12\;in}{1\;ft}\cdot\dfrac{1\;hr}{60\;min}\) That's it.

OpenStudy (yanasidlinskiy):

Hahahaha!!! lol!! Wow simple and easy!!!